what is the maximum area for a recatngle with a perimeter of 120 meters?
what is the maximum area for a recatngle with a perimeter of 130 meters?

Answer by rfer(16322)

a square gives the max area
120/4=30
30*30=900 sq meters
130/4=32.5
32.5*32.5=1065.25 sq meters